OpenStudy (narissa):

what is the scale factor of similar figures the small one is 4cm and the big one is 8.4cm

OpenStudy (narissa):

Can someone help please I'm confused with the decimal

OpenStudy (alexandervonhumboldt2):

just divide 8.4/4

OpenStudy (alexandervonhumboldt2):

8.4=84/10

OpenStudy (amorfide):

if your first shape is 4 and you want to get to the big shape which is 8.4 then your scale factor is the second shape divided by the first shape 8.4/4 is your scale factor

OpenStudy (narissa):

its 2.1 right?

OpenStudy (alexandervonhumboldt2):

84/10*1/4=84/40=2.1

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (narissa):

yeah thanks i have one more

OpenStudy (alexandervonhumboldt2):

i can help

OpenStudy (narissa):

ok this is a word problem says a dartboard is octagonal each side measures 9 in the copy of the board was smaller by a scale factor of .25 what is the side length

OpenStudy (amorfide):

alex don't help if you are not going to explain how you did it lol

OpenStudy (alexandervonhumboldt2):

just multiply 9*0.25

OpenStudy (narissa):

2.25 and all u do is multiply the side length times the smaller scale factor?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (narissa):

really that easy?

OpenStudy (triciaal):